OpenStudy (anonymous):

Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. 1. cot x sec^4x = cot x + 2 tan x + tan^3x 2. (sin x)(tan x cos x - cot x cos x) = 1 - 2 cos^2x 3. 1 + sec^2x sin^2x = sec^2x

OpenStudy (callisto):

The first one: \[LS=cot x sec^4x \]\[=\frac{cosx}{sinx}\times \frac{1}{cos^4x}\]\[=\frac{1}{sinxcos^3x}\] \[RS= cot x + 2 tan x + tan^3x\]\[=\frac{cosx}{sinx}+\frac{2sinx}{cosx}+\frac{sin^3x}{cos^3x}\]\[=\frac{cos^4x+2sin^2xcos^2x+cos^4x}{sinxcos^3x}\]\[=\frac{(cos^2x+sin^2x)^2}{sinxcos^3x}\]\[=...\]

OpenStudy (anonymous):

\[ \frac{ 1 }{\sin x \cos^3x }\]

OpenStudy (callisto):

So, you can verify the first one. Btw, do you understand my work?

OpenStudy (anonymous):

yes I understand it

OpenStudy (callisto):

Okay. So, let's move on to the next question. \[ (sin x)(tan x cos x - cot x cos x) = 1 - 2 cos^2x\] \[LS= (sin x)(tan x cos x - cot x cos x) \]\[=sinx(\frac{sinx}{cosx})cosx-sinx(\frac{cosx}{sinx})cosx\]Can you simplify it?

OpenStudy (anonymous):

I need some help

OpenStudy (callisto):

Cancel the common factors

OpenStudy (anonymous):

okay so sinx-cosx?

OpenStudy (callisto):

No. \[sinx(\frac{sinx}{cosx})cosx=sinx(\frac{sinx}{\cancel{cosx}})\cancel{cosx}=sin^2x\]

OpenStudy (anonymous):

Can you please explain it a little more?

OpenStudy (callisto):

Sure, which part?

OpenStudy (anonymous):

the simplifying

OpenStudy (callisto):

Okay. \[sinx(\frac{sinx}{cosx})cosx \]Multiply\[=\frac{(sinx)(sinx)(cosx)}{cosx}\]\[=\frac{(sin^2x)(cosx)}{cosx}\]Cancel common factor (cosx)\[=\frac{(sin^2x)\cancel{(cosx)}}{\cancel{cosx}}\]\[=sin^2x\]

OpenStudy (callisto):

For \(sinx(\frac{sinx}{cosx})cosx-sinx(\frac{cosx}{sinx})cosx\) , I've only shown you the first term. You need to work out (to simplify) the second term!

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

would it be \[\cos^2x\]

OpenStudy (callisto):

Yes!

OpenStudy (callisto):

So, \[LS= (sin x)(tan x cos x - cot x cos x)\]\[=sinx(\frac{sinx}{cosx})cosx-sinx(\frac{cosx}{sinx})cosx\]\[=sin^2x - cos^2x\]Now, rewrite \(sin^2x= 1-cos^2x\). See what you get!

OpenStudy (anonymous):

\[\sin^2x+ \cos^2x=1\]

OpenStudy (callisto):

Sorry :( I meant for \[sin^2x-cos^2x\] Replace \(sin^2x\) by \(1-cos^2x\) :S

OpenStudy (anonymous):

im a little confused

OpenStudy (callisto):

Using the identity \(sin^2x = 1-cos^2x\) \[sin^2x-cos^2x\]\[=(1-cos^2x)-cos^2x\]\[=...\]

OpenStudy (anonymous):

1-2 cos^2x?

OpenStudy (callisto):

Yes!!